Andy Mendoza makes handcrafted dolls, which he cells at craft fairs. He is considering mass-producing the dolls to sell in stores. He estimates that the initial investment for plant and equipment will be $25,000, whereas labor, material, packaging, and shipping will be about $10, per doll. If the dolls are sold for $30 each, what sales volume is necessary for Andy to break even?
Answer #1

Let the break even point be X, then

Total Expense = Total Revenue

25000 + 10*X = 30*X

20*X = 25000

X =. 1250